two cars start from points 420 miles apart and travel toward each other. They meet after 3. 5hours. Find the average speed of each car if one travels 30 mph slower than the other
Answer by mananth(16946)

Distance = 420 miles
..
speed of 1 car = x
speed of other car = x-30
...
time they meet = 3.5 hours
...
since they are travelling towards each other add up the speeds
x+x-30= 2x-30
..
t= d/r
3.5=420/(2x-30)
multiply by (2x-30)
3.5(2x-30)=420
7x-105=420
+105
7x=420+105
7x=525
/7
x=75 mph ....speed of one car
speed of other car = 75 -30 = 45 mph
...
m.ananth@hotmail.ca